OpenStudy (anonymous):

solve using the Quadratic Formula 4x^2+ 7x + 1 = 0

OpenStudy (johnweldon1993):

Well I'm assuming you know what the quadratic formula is? But here's a little catch up...Think of you coefficients as letters a, b, and c so a = 4 b = 7 and c = 1 Now use this formula \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] And find your answer...let me know how that works for you

OpenStudy (anonymous):

im confused

OpenStudy (johnweldon1993):

Which part specifically?

OpenStudy (anonymous):

putting the numbers in the quadratic formula

OpenStudy (anonymous):

\[\frac{ -7 \pm \sqrt{7^{2}-4(4)(1)} }{ 2(4) }\]

OpenStudy (johnweldon1993):

Thank you @RoseDryer , I lost connection and couldn't respond